# Online Cascade Learning for Efficient Inference over Streams

### v2

@@ Abstract
Large Language Models (LLMs) have a natural role in answering complex queries about data streams, but the high computational cost of LLM inference makes them infeasible in many such tasks. We propose online cascade learning, the first approach to address this challenge. The objective here is to learn a “cascade” of models, starting with lower-capacity models (such as logistic regression) and ending with a powerful LLM, along with a deferral policy that determines the model to be used on a given input. We formulate the task of learning cascades online as an imitation-learning problem, where smaller models are updated over time imitating the LLM expert demonstrations, and give a no-regret algorithm for the problem. Experimental results across four benchmarks show that our method parallels LLMs in accuracy while cutting down inference costs by as much as 90% with strong robustness against input distribution shifts, underscoring its efficacy and adaptability in stream processing.
@@ 2 Problem Formulation
2.1 Preliminaries
Consider a set of $N$ models, $(m_{1},...,m_{N})$, arranged in ascending order of computational cost, and an evaluation data stream $X$ consisting of user queries $x_{t}$ that arrives at time $t$. For every query $x_{t}$, the cascade system aims to output a prediction $y_{t}$ from a label set $Y$. The inference cost of a model $m_{i}$, which is assumed to be independent of $x_{t}$, is denoted as $c_{i}$. In a cascade setting, the model costs monotonically increase as $c_{1}<c_{2}<...<c_{N}$.
Therefore, we can formulate the sequential decision problem in a model cascade as an episodic Markov decision process (MDP), represented by the 4-tuple $(S,A,T,C)$:
  • States in $S$ are either a pair $\langle x_{t},i\rangle$, or a terminal state exit, where:
    • $x_{t}$ is a query from an IID set $X$ at episode $t$.
    • $i\in\{1,...,N\}$ indicates the current cascade level.
    • exit ends the current episode and resets the MDP with $\langle x_{t+1},1\rangle$ as the initial state of a new episode.
  • The set of actions $A(s)$ at any state $s$ comprises:
@@ 2.1 Preliminaries
    • A ‘defer’ action to proceed to the next level of cascade. This action is crucial in states where the model’s confidence in making an immediate prediction is low.
  • The set of deterministic transitions $T(s,a)$:
    • $\langle x_{t},i\rangle\xrightarrow[]{a\in Y}\texttt{exit}$
    • $\langle x_{t},i\rangle\xrightarrow[]{a=\texttt{defer}}\langle x_{t},i+1\rangle$
  • The cost $C(s,a)$ is defined as:
    $C(s,a)=\begin{cases}\mathcal{L}(a|x)&\text{if $a\in Y$},\\ \mu c_{i+1}&\text{if $a=${defer},}\end{cases}$
    where $\mathcal{L}(a|x)$ is the prediction loss, $c_{i+1}$ is the computational cost for the next model $m_{i+1}$, and $\mu$ is an adjustable weighting factor that guides the cost-accuracy trade-off.
2.2 Online Cascade Learning
Therefore, the online learning of a cascade is equivalent to solving the above episodic MDP, where the maximum length of each episode is $N$, the number of models in the cascade. Within this MDP, both states and actions are delineated to be independent across different episodes and input queries.
The primary objective of online cascade learning is to minimize the cumulative cost across all conceivable states over a time horizon of $T$ episodes. This cost function should encompass two critical aspects: model prediction loss and computational costs. Thus, minimizing the cumulative cost is inherently a task of achieving a delicate balance between enhancing predictive accuracy and ensuring cost efficiency.
Formally, we use $\Pi$ to denote the complete policy space under consideration, which is the theoretical aggregation of all trainable model weights in the cascade. Within this space, for a given policy $\pi$, i.e., a cascade of models of certain weights, the notation $d_{\pi}$ is employed to signify the expected state distribution when the policy $\pi$ is applied consistently throughout an entire episode. The total cost of executing policy $\pi$ for $T$ episodes can be represented as:
$J(\pi)$ $=\sum^{T}_{t=1}\mathbb{E}_{s\sim d_{\pi}}[C_{\pi}(s)],$ (1)
where $C_{\pi}(s)=\mathbb{E}_{a\sim\pi(s)}[C(s,a)]$ denotes the expected immediate cost of $\pi$ at state $s$.
Since the query stream $X$ is assumed to be IID and the states among episodes are independent, we can simplify the notation of state $\langle x_{t},i\rangle$ as $s_{i}$. Let $p(\pi,s_{i})^{\prime}$ represents the probability of choosing an action $a_{i}=\texttt{defer}$ at $s_{i}$, and $1-p(\pi,s_{i})^{\prime}$ represents the probability of choosing an action $a_{i}\in Y$. The expected immediate cost at state $s_{i}$ can be written as:
$C_{\pi}(s_{i})=(1-p(\pi,s_{i})^{\prime})\cdot\mathcal{L}(a_{i}|x_{t})+p(\pi,s_{i})^{\prime}\cdot\mu c_{i+1},$
where $\mathcal{L}(a_{i}|x_{t})$ represents the prediction loss for query $x_{t}$ if the policy chooses to output a prediction at the current state using model $m_{i}$, and $\mu c_{i+1}$ indicates the weighted computational costs to be incurred if the policy defers to the next model $m_{i+1}$ in the cascade.
Moreover, within a single episode, it is essential to accumulate the expected costs across all potential states. This necessitates accounting for the probability of a policy $\pi$ reaching a specific state $s_{i}$, denoted as $p^{s_{i}}_{\pi}$. Since a state $s_{i}$ can be reached if and only if the policy chooses to defer in all preceding states $s_{1},...,s_{i-1}$ within the current episode, therefore for a policy $\pi$, $p^{s_{i}}_{\pi}=\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}$. Please note that, the probability distribution over these states represents potential transitions, not an aggregation of model weights.
Now, the overall objective function, i.e., the cumulative cost across $T$ episodes can be expressed as:
$J(\pi)$ $=\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{i}}_{\pi}\cdot C_{\pi}(s_{i})$ (2)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}\prod^{i-1}_{j=1}p(\pi,s_{j})^{\prime}\cdot\Big{(}(1-p(\pi,s_{i})^{\prime})\cdot\mathcal{L}(a_{i}|x_{t})$
$\ \ \ \ \ +p(\pi,s_{i})^{\prime}\cdot\mu c_{i+1}\Big{)}.$ (3)
@@ 2.2 Online Cascade Learning
  Initialize $\mathcal{D}\leftarrow\emptyset$, policy weights $\hat{\pi}_{1}$, and $\beta_{1}$
  for $t=1$ to $T$ do
     Sample query $x_{t}$ from $X$
     for $i=1$ to $N$ do
        Let $\pi_{t}=\beta_{t}\pi^{*}+(1-\beta_{t})\hat{\pi}_{t}$ by setting   $\pi_{t}\leftarrow\pi^{*}$ at a decaying probability $\beta_{t}$
        $a^{i}_{t}=\pi_{t}(s_{i})$
        Continue if $a^{i}_{t}=\texttt{defer}$
        $y_{t}\leftarrow a^{i}_{t}$
        $\mathcal{D}\leftarrow\mathcal{D}\cup\{x_{t},y_{t}\}$ if $i$ reaches $N$ or $\pi_{t}=\pi^{*}$
@@
     Output $y_{t}$
     Compute $J(\pi_{t})$ on $\mathcal{D}$ and update $\hat{\pi}_{t+1}$ via OGD
@@
Therefore, by collecting the annotations from an expert policy as the approximation of ground truth labels, we can optimize both models (i.e., $m_{1},...,m_{N}$) and deferral policies (i.e., $p(\pi,s_{1})^{\prime},...,p(\pi,s_{N-1})^{\prime}$) together by minimizing the above objective function via online gradient descent.
3 Methodology
We now present our proposed online cascade learning algorithm that updates the policy by imitating the behaviors of an expert policy (i.e., LLM) like DAgger (Ross et al., 2011) to balance cost efficiency and good performance guarantee.
The overall process of the algorithm is detailed in Algorithm 1, where $\pi^{*}$ refers to the LLM expert for imitation, and $\hat{\pi}_{1}$ indicates the initial states of the small models. In practice, the models are initialized either randomly or with their respective pretrained weights. For every incoming query $x_{t}$ at episode $t$ (outer loop), we sequentially utilize the $i$-th model in the current cascade to make a prediction (inner loop). Specifically, the model outputs a class probability distribution, which is converted by the policy into a deferral probability score to determine the action $a^{i}_{t}=\pi_{t}(s_{i})$
If the action is defer, the query would be navigated to the cascade’s next $(i+1)$-th model. Otherwise, if the action is to output a prediction at the current cascade level, it will make a prediction $y_{t}\leftarrow a^{i}_{t}$, and break the inner loop (succeeding models will not be activated for current query $x_{t}$). If the $N$-th model in the cascade is reached, that means the query has been deferred to the LLM expert $\pi^{*}$, thereby we collect its annotation $\{x_{t},y_{t}\}$ and aggregate it to the dataset $\mathcal{D}$.
By default, $\pi_{t}(s_{i})$ invokes the $i$-th model $\hat{\pi}_{t}(s_{i})=m_{i}^{t}(x_{t})$ for making an action at episode $t$. However, throughout the decision-making process, the cascade may optionally skip the rest layers and directly jump to the LLM expert $\pi^{*}$ at a non-zero decaying probability $\beta_{t}$ to obtain its demonstrations $\{x_{t},y_{t}\}$, similar to DAgger (Ross et al., 2011), for faster convergence. Therefore, in Algorithm 1, $\pi_{t}=\beta_{t}\pi^{*}+(1-\beta_{t})\hat{\pi}_{t}$ represents an expected value of the acting policy, not the combination of model parameters.
After processing each query, it outputs the prediction $y_{t}$, then update the policy $\hat{\pi}_{t+1}$, including the small models and deferral policy, to mimic the LLM expert on the collected trajectories $\mathcal{D}$. At a high level, the algorithm continuously collects annotations from the expert (e.g., at a decaying probability $\beta_{t}$ or when the state reaches $s_{N}$ in the episode) and updates the policy through online gradient descent based on expert demonstrations. Practically, the user can change the cost weighting factor $\mu$ in the objective function $J(\pi)$ and the initial decaying factor $\beta_{1}$ for adjusting cost budgets.
3.1 Theoretical Analysis
In online learning, a policy’s regret over time $T$ is its total cost minus the cost of the best fixed policy in hindsight. For online cascade learning, the regret of a learned policy $\pi$ can be defined as:
$\gamma$ $=J(\pi)-\min\limits_{\pi\in\Pi}J(\pi)$ (4)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{i}}_{\pi}C_{\pi}(s_{i})-\min\limits_{\pi\in\Pi}\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{i}}_{\pi}C_{\pi}(s_{i})$ (5)
We begin our analysis by examining a simplified online ensemble learning algorithm, comprising a series of models $m_{1},...,m_{N}$, each with a fixed operating probability $\sum_{i=1}^{N}w_{i}=1$. Assuming a convex, differentiable cost function $c^{t}$ for all $t$, and bounded, closed, nonempty model spaces $||M_{i}||$ for all $i$, we analyze the regret of this algorithm.
@@ 3.1 Theoretical Analysis
For online cascade learning with $\eta_{t}=t^{-1/2}$, the algorithm’s total regret is $o(T)$, implying $\lim_{T\rightarrow\infty}\gamma/T\leq 0$, i.e., the average regret approaches zero as $T$ grows.
Having established the no-regret property for online ensemble learning, we extend this to our online cascade learning algorithm. By replacing the fixed probabilities $w_{i}$ in Equation (3.1) with dynamic probabilities based on preceding model actions, and demonstrating convergence of these dynamic probabilities to optimal values as $T\rightarrow\infty$ (proof in Appendix A, Lemma A.2), we mirror the total regret of online cascade learning to that of the online ensemble learning algorithm, completing the proof for Theorem 3.2. Further details are in Appendix A.
3.2 Confidence Calibration
A reliable deferral policy is crucial for a cascade system to determine whether to invoke the next model (i.e., to defer) or to output the predictions. Currently, most existing works make deferral decisions based on a confidence score, typically measured either by the maximum predictive probability across all classes (Wang et al., 2022; Varshney & Baral, 2022), or the Shannon entropy of the predictive distribution (Stogiannidis et al., 2023).
However, for online cascade learning, where the model capabilities are dynamically updated, and the annotations from LLMs may be noisy, confidence-based deferral policies have been shown to be inadequate (Jitkrittum et al., 2023). To aid the learning of deferral policies (i.e., $p(\pi,s_{i})^{\prime}$), we adopt a post-hoc deferral policy $\pi_{i}^{\prime}$ to calibrate the confidence estimate of a prediction from model $m_{i}$. It is implemented using a multi-layer perceptron (MLP) that takes the corresponding model’s predictive probabilities as input, and the policy can be learned with the following objective:
$\min_{\pi_{i}^{\prime}:\mathbb{R}^{|Y|}\rightarrow\mathbb{R}}\sum_{x_{t}\in X}\sum_{i=1}^{N-1}L(\pi^{\prime}_{i}(m_{i}(x_{t})),z_{i}),$ (7)
@@ 3.2 Confidence Calibration
In our work, since inference is performed in a zero-shot setting without any human annotations, we use the predictions of the LLM (i.e., $\operatorname*{arg\,max}m_{N}(x_{t})$) as the approximation of the ground truth labels $y^{*}_{t}$. Thus, calibration is only performed on those input queries where the LLM is invoked. During the calibration, the learning of $\pi_{i}^{\prime}$ would only optimize the parameters of the MLP, not the models $m_{i}$ in the cascade. During inference time, the post-hoc deferral policy would choose to defer if $\pi^{\prime}_{i}(m_{i}(x_{t}))>0.5$. Otherwise, it would output the prediction of the current model.

### v3

@@ Abstract
Large Language Models (LLMs) have a natural role in answering complex queries about data streams, but the high computational cost of LLM inference makes them infeasible in many such tasks. We propose online cascade learning as an approach to address this challenge. The objective here is to learn a “cascade” of models, starting with lower-capacity models (such as logistic regression) and ending with a powerful LLM, along with a deferral policy that determines the model to be used on a given input. We formulate the task of learning cascades online as an imitation-learning problem, where smaller models are updated over time imitating the LLM expert demonstrations, and give a no-regret algorithm for the problem. Experimental results across four benchmarks show that our method parallels LLMs in accuracy while cutting down inference costs by as much as 90% with strong robustness against input distribution shifts, underscoring its efficacy and adaptability in stream processing.
@@ 2 Problem Formulation
Inference over Streams as an MDP Problem
We consider stream processing scenarios that have as input a fixed infinite stream $X=\langle x_{1},\dots,x_{t},\dots\rangle$ of user queries. The $t$-th query $x_{t}$ is associated with a ground-truth label $y_{t}\in Y$, where $Y$ is a label set. Our goal is to predict the label for each $x_{t}$ using an $N$-level model cascade.
We formulate our problem using an episodic Markov decision process (MDP) $(S,A,\mathcal{T},C)$. Here:
  • $S$ is a set of states. A state in the $t$-th episode is either a pair $\langle x_{t},i\rangle$, where $i\in\{1,...,N\}$ indicates the current cascade level, or a special terminal state exit that ends the episode. The initial state of the $t$-th episode is $\langle x_{t},1\rangle.$ For clarity, we abbreviate $\langle x_{t},i\rangle$ by $s_{t,i}$.
  • $A$ is a set of actions, consisting of:
@@
    • A special action defer that activates the next level of the cascade.
  • $\mathcal{T}(s_{t,i},a)$ is a deterministic transition function, consisting of transitions of the form:
    • $\mathcal{T}(s_{t,i},a)=\texttt{exit}$ for $a\in Y$.
    • $\mathcal{T}(s_{t,i},\texttt{defer})=s_{t,i+1}$.
  • $C(s_{t,i},a)$ is a cost function defined as:
    $C(s_{t,i},a)=\begin{cases}\mathcal{L}(a|y_{t})&\text{if $a\in Y$},\\ \mu c_{i+1}&\text{if $a=${defer}.}\end{cases}$
    Here, $\mathcal{L}(a|y_{t})$ is a prediction loss that measures the accuracy of the cascade’s prediction. $c_{i+1}$ represents the penalty we pay for a deferral — intuitively, this penalty captures the computational overheads of going one level deeper into the cascade. The adjustable constant $\mu$ guides the trade-off between computational cost and accuracy.
Online Cascade Learning
We now formulate online cascade learning as the problem of solving the above episodic MDP. Let a policy $\pi$ be a stochastic map from states to actions. We use $\pi(s_{t,i},\texttt{defer})$ to represent the probability that $\pi$ chooses to defer in state $s_{t,i}$, and $\pi(s_{t,i},y)$ to represent the probability that $\pi$ chooses $y\in Y$, conditioned on no deferral having occurred.
The probability of a policy $\pi$ entering state $s_{t,i}$ in episode $t$ is denoted as $p^{s_{t,i}}_{\pi}$. For $i>1$, $s_{t,i}$ can be reached if and only if the policy chooses the defer action in all preceding states $s_{t,1},...,s_{t,i-1}$ within the current episode. Thus,
$p^{s_{t,1}}_{\pi}=1,\qquad\qquad p^{s_{t,i}}_{\pi}=\prod^{i-1}_{j=1}\pi(s_{t,j},\texttt{defer}).$
Then, the cost of executing $\pi$ over $T$ episodes is computed by summing over all the episodes and cascade levels:
$J(\pi,T)=\sum_{t=1}^{T}\left[\sum_{i=1}^{N}p^{s_{t,i}}_{\pi}C_{\pi}(s_{t,i})\right]$ (1)
Here, $C_{\pi}(s_{t,i})$ is the expected, immediate cost of applying policy $\pi$ at state $s_{t,i}$. This is computed as:
$C_{\pi}(s_{t,i})$ $=\pi(s_{t,i},\texttt{defer})\cdot\mu c_{i+1}$
$+(1-\pi(s_{t,i},\texttt{defer}))\cdot\sum_{y\in Y}\pi(s_{t,i},y)\cdot\mathcal{L}(y|y_{t}).$
After having seen the first $T$ queries in the input stream $X$, our learning goal is to find a policy that minimizes $J(\pi,T)$. When $T$ is clear from the context, we often abbreviate $J(\pi,T)$ by $J(\pi)$.
Policy Representations
We represent policies in a factorized way using a set of classification models $\langle m_{1},\dots,m_{N}\rangle$ that constitute the different levels of the cascade, and a set of deferral functions $\langle f_{1},\dots,f_{N-1}\rangle$ that decide whether the current level can perform a high-confidence classification or to defer. We assume each $m_{i}$ to produce a vector of probabilities, with one probability for each label, and each $f_{i}$ to produce a probability of deferral. Then the overall policy has the form:
$\pi(s_{t,i},\texttt{defer})$ $=f_{i}(m_{i}(x_{t})),$
$\pi(s_{t,i},y)$ $=\left(m_{i}(x_{t})\right)[y]\textrm{ for }y\in Y.$
We assume the two parameterized function representations for each level $i$: the classification model $m_{i}$ and the deferral function $f_{i}$, are both characterized by a crucial property: they guarantee uniform computational costs for evaluating the functions, regardless of the specific function parameters and function inputs. This means that for any instantiation of the parameters in $m_{i}$ and $f_{i}$, the inference costs remain constant, irrespective of the specific input query. This assumption is reasonable in the practical scenarios we target. For example, the inference cost of a BERT-base model is approximately the same, no matter how it is parameterized. This uniform cost assumption also underpins the use of fixed cost penalties $c_{i}$ in our MDPs.
3 Learning Algorithm
To learn cascades online without human annotations, we propose an imitation learning algorithm that assumes an expert policy (an LLM) that can demonstrate ground-truth labels. The algorithm iteratively updates the classification models and deferral functions in the cascade by imitating the expert as in DAgger (Ross et al., 2011). However, unlike traditional imitation learning, our goal here is to balance computational efficiency and accuracy.
We incorporate such an expert into our cascades by assuming that the final classification model $m_{N}$ is the expert LLM. When invoked on a query $x_{t}$, $m_{N}$ always outputs a vector whose largest value is associated with ground-truth label $y_{t}$. However, it may not always be the case that choosing to invoke $m_{N}$ leads to optimal cost. There may be a smaller model $m_{i}$ for $i<N$ that also produces the correct label $y_{t}$, without the cost of the additional defer actions. Indeed, an optimal policy may occasionally incur prediction errors to manage the cost of incurring defer actions. Designing an algorithm that can train a policy in an online fashion to manage those trade-offs is at the core of the paper.
Our overall algorithm is detailed in Algorithm 1. Here, $m_{1},...,m_{N}$ are the classification models (with $m_{N}$ being the expert for imitation) and $f_{1},...,f_{N-1}$ are the deferral functions. The smaller models $m_{1},...,m_{N-1}$ are initialized either randomly or with their respective pretrained weights. For each incoming query $x_{t}$ (outer loop), we sequentially utilize the $i$-th model in the cascade to make a prediction (inner loop). Specifically, the model generates a probability vector $\textit{pred}_{i}$, which is processed by the deferral function $f_{i}$. This yields a deferral probability score that can be discretized into a binary $\textit{action}_{i}$ to determine whether the prediction at the current cascade level is reliable.
If the action is defer, the query would be navigated to the cascade’s next $(i+1)$-th model. Otherwise, it will make a prediction $\hat{y}_{t}=\operatorname*{arg\,max}(\textit{pred}_{i})$, and break the inner loop (succeeding models will not be activated for current query $x_{t}$). If the query has been deferred to the LLM expert $m_{N}$ at the last cascade level, its annotation $\hat{y}_{t}$ is regarded as the ground truth $y_{t}$ and aggregated to the dataset $\mathcal{D}$.
Throughout the inner iteration, at cascade level $i$, it may optionally skip the rest layers and jump to the LLM expert $m_{N}$ at a non-zero decaying probability $\beta_{i}$ to directly obtain its demonstration and aggregate it to dataset $\mathcal{D}$, similar to DAgger (Ross et al., 2011), for faster convergence.
@@ 3 Learning Algorithm
  Initialize models $m_{1},...,m_{N}$, deferral functions $f_{1},...,f_{N-1}$, $\beta_{1}$, $\mathcal{D}\leftarrow\emptyset$
  for $x_{t}$ in stream $X$ do
     for $m_{i}$ in $m_{1}$ to $m_{N}$ do
        At probability $\beta_{t}$:
            Jump to $m_{N}$ // like DAgger
        $\textit{pred}_{i}=m_{i}(x_{t})$ // probability vector
        $\textit{action}_{i}\sim f_{i}(\textit{pred}_{i})$ // whether to defer
        if $m_{i}$ is $m_{N}$ or $\textit{action}_{i}\neq\texttt{defer}$ then
           $\hat{y}_{t}=\operatorname*{arg\,max}(\textit{pred}_{i})$ // discretization
           $\mathcal{D}\leftarrow\mathcal{D}\cup\{x_{t},\hat{y}_{t}\}$ if $m_{i}$ is $m_{N}$
@@
     Output $\hat{y}_{t}$
     Update $m_{1}$ to $m_{N-1}$ on $\mathcal{D}$ via OGD // imitating expert
     Compute loss $J(\pi,t)$ and update $f_{1}$ to $f_{N-1}$ via OGD
     Decay $\beta_{t+1}$
@@
After processing each query, the algorithm outputs the prediction $y_{t}$, then updates the small models $m_{1},...,m_{N-1}$ to mimic the expert demonstrations on the collected trajectories $\mathcal{D}$. Similarly, the deferral functions $f_{1},...,f_{N-1}$ are also updated based on the loss computed by Equation (1). The algorithm continuously collects annotations from the LLM expert (e.g., at a decaying probability $\beta_{t}$ or when the query is deferred to $m_{N}$) and updates the policy via online gradient descent (OGD). Practically, the user can change the cost weighting factor $\mu$ in the loss function $J(\pi)$ and the initial decaying factor $\beta_{1}$ for adjusting cost budgets.
Theoretical Analysis
In online learning, a policy’s regret over time $T$ is its total cost minus the cost of the best fixed policy in hindsight. In our setting, the regret of a learned policy $\pi$ is:
$\gamma$ $=J(\pi,T)-\min\limits_{\pi\in\Pi}J(\pi,T)$ (2)
$=\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{t,i}}_{\pi}C_{\pi}(s_{t,i})-\min\limits_{\pi\in\Pi}\sum^{T}_{t=1}\sum^{N}_{i=1}p^{s_{t,i}}_{\pi}C_{\pi}(s_{t,i}),$ (3)
where $\Pi$ denotes the whole possible policy space. An algorithm is defined to have no-regret if it can produce a sequence of policies $\pi_{1},\dots,\pi_{T}$ such that the expected average regret $\gamma/T$ goes to 00 as $T\rightarrow\infty$ (Ross et al., 2011).
To aid our no-regret analysis of online cascade learning, we start by constructing a simplified online ensemble learning algorithm under the same stream processing setting that comprises the linear combination of a series of classification models $m_{1},...,m_{N}$, each with a static operating probability $\sum_{i=1}^{N}w_{i}=1$, without any deferral functions. Let us denote the model parameters of $m_{i}$ at time $t$ by $m_{i}^{t}$. Assuming a convex, differentiable cost function $c^{t}$ for all $t$ that can evaluate $m_{i}^{t}$, and bounded, closed, nonempty model spaces $||M_{i}||$ for all $m_{i}$, we analyze the regret of this algorithm.
@@
For online cascade learning with $\eta_{t}=t^{-1/2}$, the algorithm’s total regret is $o(T)$, implying $\lim_{T\rightarrow\infty}\gamma/T\leq 0$, i.e., the average regret approaches zero as $T$ grows $\infty$.
Having established the no-regret property for online ensemble learning, we extend this to our online cascade learning algorithm. By replacing the static probabilities $w_{i}$ in Equation (3.1) with dynamic probabilities based on preceding model actions, and demonstrating convergence of these dynamic probabilities to optimal values as $T\rightarrow\infty$ (proof in Appendix A, Lemma A.2), we mirror the total regret of online cascade learning to that of the online ensemble learning algorithm, completing the proof for Theorem 3.2. Further details are in Appendix A.
Confidence Calibration
A reliable deferral rule is crucial for a cascade system to determine whether to invoke the next model (i.e., to defer) or to output the predictions. Currently, most existing works make deferral decisions based on a confidence score, typically measured either by the maximum predictive probability across all classes (Wang et al., 2022; Varshney & Baral, 2022), or the Shannon entropy of the predictive distribution (Stogiannidis et al., 2023).
However, for online cascade learning, where the model capabilities are dynamically updated, and the annotations from LLMs may be noisy, confidence-based deferral rules have been shown to be inadequate (Jitkrittum et al., 2023). To aid the learning of deferral functions (i.e., $f_{1},...,f_{N-1}$), we adopt a post-hoc approach to calibrate the confidence estimate of a certain model’s prediction $m_{i}(x_{t})$. It is implemented using a multi-layer perceptron (MLP) that takes the corresponding model’s predictive probabilities as input, and the functions can be updated with the following objective:
$\min_{\pi_{i}^{\prime}:\mathbb{R}^{|Y|}\rightarrow\mathbb{R}}\sum_{x_{t}\in X}\sum_{i=1}^{N-1}L(f_{i}(m_{i}(x_{t})),z_{i}),$ (5)
@@
Since we treat the expert LLM predictions (i.e., $\operatorname*{arg\,max}(m_{N}(x_{t}))$) as the ground truth labels $y_{t}$, calibration is only performed on those input queries where the expert LLM is invoked. During the calibration, the learning of $f_{i}$ would only optimize the parameters of the MLP, not the models $m_{i}$ in the cascade. During inference time, the post-hoc deferral function would choose to defer if $f_{i}(m_{i}(x_{t}))>0.5$. Otherwise, it would output the prediction of the current model.